OpenStudy (anonymous):

Identify the relative minimum value for the function shown below. g(x)=x^3-3x^2+2

OpenStudy (anonymous):

(2,-2) i think

OpenStudy (anonymous):

g(x)=x^3-3x^2+2 g'(x)=3x^2-6x=3x(x-2) x=0 or x=2 plug in both values and u get that the minimum should be (2,-2)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw